Titration of a diprotic acid with a strong base You have a 10.0 mL solution containing 0.5 M carbonic acid. Carbonic acid is diprotic, with pKa1 = 6.35 and pKa2 = 10.33. You titrate this solution using 1.00 M NaOH .

(a) Calculate the pH of the solution before adding any NaOH.

(b) Calculate the amount of NaOH needed to reach the first midpoint. What is the pH?

(c) Calculate the amount of NaOH needed to reach the first equivalence point. What is the pH?

(d) Calculate the amount of NaOH needed to reach the second midpoint. What is the pH?

(e) Calculate the amount of NaOH needed to reach the second equivalence point. What is the pH?

(f) Calculate the pH of the solution after adding 15 mL of NaOH.

(g) Use your above calculations to sketch the pH for the solution during the entire titration.
Answer #1

Ž HCO3 + H+ Ango - H₂CO3 prag = 6.35. - KO . Ka = 4.407 107 HCO3 co²- + Ht i n - pka = 10.33 Kl = 4.67 x 10 Let H₂ W3 = H₂A(6 At mild of pt. equivalance point. [H₂A) = [Na] pH = pka, = 6.35 u volume of IM NaOH is needed M,V, = M2 V2 1 = 0.5 x 10 ml1x V ml = V 0.5x 10 = 5 ml (d) At point mid of and equivalance - THAT = [42 ] pH = pka, pH = 10.33 Let Ve volume of IM NaOH iFage no 42- + 0.25 H₂O - & HA o OH one 0.25-x k = 2.14x10 4 - x² O 0.25-2 xca 0.25 x= 0.25 x 2.14x10-4 x = 0.73x102 N = 7.3